# Topological frequency conversion in rhombohedral multilayer graphene

### v1

@@ Abstract
Quantum systems coupled to light exhibit a range of unique emergent phenomena, providing a powerful platform to design and control phases of matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. It was recently appreciated that the paradigm of Floquet engineering can be generalized to multi-chromatic driving, leading to even richer physics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 31]. In this setting, the photon numbers of distinct modes act as synthetic lattice coordinates, opening access to higher-dimensional topological phases that support novel pumping mechanisms. A prime example is topological frequency conversion, whereby energy is transferred between two modes with frequencies $f_{1}$ and $f_{2}$ at a rate given by the power quantum $hf_{1}f_{2}$. In the solid-state realm, Weyl semimetals were recently proposed as a platform to realize this effect, leveraging the sources and sinks of Berry curvature carried by their Weyl points [31].
@@
Here we propose rhombohedral multilayer graphene (RMG) [33, 34] as an alternative platform for realizing topological frequency conversion. Similarly to Weyl semimetals, RMG hosts Weyl nodes, but in a three-dimensional hybrid space formed by the two crystal momenta and the interlayer potential difference. This hybrid space can be dynamically addressed through the in-plane vector potential and out-of-plane electric field induced by the driving modes. Under suitable configurations of incident beams, a fraction of charge carriers in the system converts energy between the modes at the quantized rate $hf_{1}f_{2}$, as illustrated in Fig. 1.
@@
Figure 1: (a) Rhombohedral multilayer graphene can mediate topological frequency conversion between two incident driving modes with frequencies $f_{1}$, $f_{2}$. An interlayer potential $u$ gaps out the low-energy bands of RMG [shown for bilayers in (b) and trilayers in (c)], with a Berry curvature controlled by ${\rm sgn}(u)$. (d) The driving modes couple to electrons via their in-plane vector potential and perpendicular electric field, contained in the 3d vector $\bm{\alpha}$ [Eq. 5]. The Berry curvature ${\bm{\Omega}}$ in $\bm{\alpha}$ space (black arrows) hosts a monopole or Weyl node located at $(-k^{x},-k^{y},0)$ for an electron with wavevector $(k^{x},k^{y})$. Under quasi-periodic driving, the system explores a 2d surface ${\cal S}_{\bm{\alpha}}$, shown for the drive configuration in (a). Frequency conversion power is given by the quantized flux of Berry curvature $\mathbf{\Omega}$ through ${\cal S}_{\bm{\alpha}}$. (e) The intersection between ${\cal S}_{\bm{\alpha}}$ and the $E^{z}=0$ plane defines a momentum area $\cal A_{\bm{\alpha}}$ where electrons contribute to conversion power at a quantized rate $hf_{1}f_{2}L$. Contours are labeled by the ratio of electric fields $E^{z}_{0}/E_{1}$.
The quasi two-dimensional nature of RMG offers key advantages including gate tunability, reduced screening and lack of skin effects. Moreover, the different type of coupling to in-plane and out-of-plane radiation fields leads to a strong sensitivity to the angle and polarization of the incident beams. For THz implementations, Bernal bilayer graphene appears most promising due to its pristine nature and lower dissipation. We finally comment on strategies to circumvent cancellations between the two (time-reversed) valleys of graphene by inducing valley-asymmetric electronic distributions.
@@
Setup. RMG consists of $L$ graphene sheets stacked such that each $A$ sublattice lies above the $B$ sublattice of the layer below—these superimposed sites are coupled with the leading interlayer tunneling term $\gamma_{1}\approx 360$ meV. Near charge neutrality, the wavefunctions of carriers primarily reside in the subspace ($A_{1},B_{L}$) comprising the leftover $A$ sites of the bottom layer and $B$ sites of the top layer. The system can thus be described by an effective two-level Hamiltonian in each of the two valleys [34],
$H_{\xi}(\mathbf{k})=\begin{pmatrix}u/2&-\gamma_{L}\left(\xi k_{x}-ik_{y}\right)^{L}\\ -\gamma_{L}\left(\xi k_{x}+ik_{y}\right)^{L}&-u/2\end{pmatrix},$ (3)
@@
with the wavevector $\mathbf{k}$ measured relative to the hexagonal Brillouin zone (BZ) corner corresponding to valley $\xi=\pm 1$. Here $u$ is an interlayer potential difference and $\gamma_{L}=(\hbar v_{F})^{L}/\gamma_{1}^{L-1}$ where $v_{\rm F}=8.4\times 10^{5}$ m/s denotes the Fermi velocity of graphene. For $u=0$ the low-energy excitations of RMG are chiral fermions with an $L$-th power dispersion [35]. Applying an out-of-plane electric field $E^{z}$ generates a potential $u=d(L-1)eE^{z}$, with $d=0.33$ nm the interlayer distance, opening a gap in the dispersion and controlling the Berry curvature of the low-energy bands (Fig. 1b)—a feature at the heart of our proposal.
@@
Bichromatic driving induces an electric field ${\bf{E}}(t)={\bf{E}}_{1}(t)+{\bf{E}}_{2}(t+\Delta t)$ in the system, where ${\bf{E}}_{j}(t)$ is the monochromatic field from mode $j$ with frequency $f_{j}$, and $\Delta t$ controls the relative phase shift between the modes—relevant when $f_{1}$ and $f_{2}$ are commensurate. Due to the quasi 2D nature of RMG we disregard skin effects and plasmon excitations that become important in 3D implementations such as Weyl semimetals [31]. We further consider THz-scale frequencies, corresponding to wavelengths $\lambda\sim 100\,\rm\mu m$, such that the electric field can be considered uniform in the sample. The in-plane component of the ${\bf{E}}$, ${\bf{E}}^{\parallel}$, couples to electronic crystal momenta via the Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}+e\mathbf{A}^{\parallel}(t)/\hbar$, with ${\bf{A}}^{\parallel}(t)$ the net vector potential induced by ${\bf{E}}^{\parallel}(t)$. In contrast, the out-of-plane component of ${\bf{E}}$, $E^{z}$, couples through the effective dipole moment in the low-energy subspace, Eq. (3). The coupling of radiation thus strongly depends on the incidence angles $\theta_{j}$ and polarization of the drives (Fig. 1a), a unique feature of this setup.
@@
${\bf{d}}_{\xi}(\mathbf{k},t)=\gamma_{L}\left(\text{Re}[\Pi_{\xi}^{L}(\mathbf{k},t)],\text{Im}[\Pi_{\xi}^{L}(\mathbf{k},t)],\alpha^{z}(t)\right).$ (4)
@@
Below, we find it convenient to re-express $\bm{\alpha}_{j}$ (and therefore $\mathbf{d}_{\xi}$) explicitly in terms of the phases $\phi_{j}=\omega_{j}t$, with $\omega_{j}=2\pi f_{j}$ the angular frequency of mode $j$.
@@
When the time dependence of $H_{\xi}(\mathbf{k},t)$ is quasi-adiabatic, $P_{\rm FC}^{\xi}(\mathbf{k})$ is proportional to the number of times the surface spanned by the unit vector $\hat{\mathbf{d}}_{\xi}\equiv{\bf{\mathbf{d}_{\xi}}}/|{\bf{\mathbf{d}_{\xi}}}|$ encloses the origin, $P_{\rm FC}^{\xi}(\mathbf{k})=hf_{1}f_{2}\int_{0}^{2\pi}\frac{d\phi_{1}d\phi_{2}}{4\pi}\hat{{\bf{d}}}_{\xi}\cdot(\partial_{\phi_{1}}\hat{{\bf{d}}}_{\xi}\times\partial_{\phi_{2}}\hat{{\bf{d}}}_{\xi})$ [17, 31, 36]. This result is derived in the Appendix and emerges from the leading-order diabatic correction to the dynamics—along with the fact that, for incommensurate frequencies, the system uniformly explores the two-dimensional parameter space $0\leq\phi_{j}\leq 2\pi$. For commensurate frequencies, this formula is recovered by also averaging $P_{\rm FC}^{\xi}(\mathbf{k})$ over $\Delta t$. Changing variables from $\phi_{j}$ to $\bm{\alpha}$ yields
@@
In analogy to Weyl semimetals, sources and sinks of Berry curvature in the hybrid three-dimensional space $\bm{\alpha}$ are quantized and confined to topologically-protected gap-closing points, analogous to Weyl nodes  [^1] The term Weyl node normally refers to such nodal points in the parameter space comprising the crystal momenta in 3D reciprocal space. Here we use the term “Weyl node” in a generalized sense to describe gap closings as a function of any three parameters.. Indeed, the RMG Hamiltonian in Eq. (4) hosts a charge-$L$ Weyl node at $\bm{\alpha}_{0}(\mathbf{k})=(-k^{x},-k^{y},0)$,
$\nabla_{\bm{\alpha}}\cdot{\mathbf{\Omega}_{\xi}}(\mathbf{k},\bm{\alpha})=-2\pi\xi L\delta(\bm{\alpha}-\bm{\alpha}_{0}(\mathbf{k})).$ (7)
Electrons therefore become topological frequency converters—mediating an energy transfer at the quantized rate $P_{\rm FC}^{\xi}(\mathbf{k})=-hf_{1}f_{2}\xi L$—if located within the BZ area delineated by the intersection of the surface $\cal S_{\bm{\alpha}}$ and the $\alpha^{z}=0$ plane of the Weyl node, denoted $\cal A_{\bm{\alpha}}$ (see Fig. 1d, e). For electrons outside of $\cal A_{\bm{\alpha}}$, $P_{\rm FC}^{\xi}(\mathbf{k})=0$.
We now consider the dynamics of the many-body state. In the regime of adiabatic driving and slow relaxation relative to the driving frequency, the steady-state population of band $n=0,1$ near valley $\xi$ takes an approximately stationary value, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$  [^2] In a simple relaxation-time approximation, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$ is given by the time-averaged value of $f_{\beta}([\varepsilon_{n}(\mathbf{k},t)-\mu])$ with $f_{\beta}(E)$ the Fermi-Dirac distribution [31]; we expect more genral dissipation mechanisms will lead to qualitatively similar distributions.. The frequency conversion intensity (i.e. the conversion power per unit area) from valley $\xi$ is obtained by weighing the individual contribution $P_{\rm FC}^{\xi}(\mathbf{k})$ of each electron with the corresponding steady-state occupation [31],
${\cal I}_{\rm FC}^{\xi}=hf_{1}f_{2}L\xi\int_{\mathcal{A}_{\alpha}}\frac{{\rm d}^{2}\mathbf{k}}{(2\pi)^{2}}\left(\bar{\rho}_{1}^{\xi}(\mathbf{k})-\bar{\rho}_{0}^{\xi}(\mathbf{k})\right).$ (8)
The boundary of $\mathcal{A}_{\bm{\alpha}}$ contains the wavevectors $\mathbf{k}$ for which the instantaneous gap of $H_{\xi}(\mathbf{k},t)$ closes at some time $t$. States near this boundary thus undergo inter-band Landau-Zener tunneling, which causes significant dissipation through Landau-Zener absorption—whereby electrons pumped to the valence band relax dissipatively, leading to an irreversible loss of the invested energy [31].
@@
[…]
When the chemical potential $\mu_{\xi}$ in a given valley moves away from charge neutrality, we expect the conversion rate $I_{\rm FC}^{\xi}$ to decrease, due to the cancellation between conduction and valence-band carriers (for $\mu_{\xi}>0$) or de-populating the valence band (for $\mu_{\xi}<0$). Because of the opposite contributions from carriers in valleys $\xi=\pm 1$ (a consequence of time-reversal symmetry), a net effect requires a valley imbalance in the number of active frequency converters, either through unequal chemical potentials in each valley, $\mu_{+}\neq\mu_{-}$, or a valley-selective population inversion. We discuss various strategies to induce such imbalances below and in the Appendix.
Numerical simulations. We support our conclusions with numerical simulations of Bernal bilayer graphene subjected to bichromatic radiation, including trigonal warping corrections to the low-energy band structure [39]. Following Ref. [31], we model the system through a master equation
$\partial_{t}\hat{\rho}_{\xi}(\mathbf{k},t)=-\frac{i}{\hbar}[\hat{H}_{\xi}(\mathbf{k},t),\hat{\rho}_{\xi}(\mathbf{k},t)]+{\cal D}_{\xi}(\mathbf{k},t)\circ\hat{\rho}_{\xi}(\mathbf{k},t),$ (10)
where $\hat{H}_{\xi}(\mathbf{k},t)$ and $\hat{\rho}_{\xi}(\mathbf{k},t)$ denote the second-quantized Hamiltonian and density matrix of the system in the Fock space spanned by electronic modes with wavevector $\mathbf{k}$ in valley $\xi$. We model dissipation through ${\cal D}_{\xi}(\mathbf{k},t)=-\frac{1}{\tau}(\hat{\rho}_{\xi}(\mathbf{k},t)-\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t))$, characterizing uniform relaxation on a timescale $\tau$ towards the instantaneous equilibrium state $\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t)=e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}/{\rm Tr}[e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}]$, with $\beta$ the inverse temperature of the bath and $\hat{n}$ the density operator. For simplicity we work at charge neutrality, $\mu_{\xi}=0$; we consider effects of driving in the presence of a Fermi surface (for $\mu_{\xi}\neq 0$) in the Appendix.
Figure 2: Momentum-resolved frequency conversion in Bernal bilayer graphene. We consider two linearly-polarized modes in the configuration of Fig. 1a with $\left(\theta_{j}-\pi/2\right)\approx 0.1^{\circ}$, relaxation time $\tau=500$ ps and charge neutrality, $\mu=0$. Contributions to frequency conversion from valleys $\xi=+1$ (a,d) and $-1$ (b,e) are shown alongside dissipated power (c,f). Frequencies $f_{1}\approx 0.128$ THz and $f_{2}=f_{1}/\sqrt{2}$ [(a)–(c)] or $f_{2}=2f_{1}/3$ [(d)–(f)] illustrate quasi-periodic and periodic driving. The color scales indicate time-averaged conversion power in units of $hf_{1}f_{2}$. The overlaid contour shows the boundary of ${\cal A}_{\alpha}$.
@@
This expression can be derived from first principles and, being agnostic to our theoretical analysis, serves as an independent check of its conclusions.
@@
We consider two linearly-polarized modes propagating with incidence angles $\theta_{j}$ (relative to $\hat{\mathbf{z}}$) in the $xz$ and $yz$ planes, as depicted in Fig. 1a. The induced vector potentials are given by $\mathbf{A}_{1}(t)=A_{1}(\cos\theta_{1}\cos[\omega_{1}t],0,\sin\theta_{1}\cos[\omega_{1}t])$ and $\mathbf{A}_{2}(t)=A_{2}(0,-\cos\theta_{2}\cos[\omega_{2}(t{+\Delta t})],\sin\theta_{2}\cos[\omega_{2}(t{+\Delta t})])$, with $\Delta t$ controlling the phase shift between the modes. We consider large angles of incidence $\theta\approx\pi/2$, to compensate for the weaker effective coupling between the radiation fields and $\sigma^{z}$ in Eq. 3. In order to obtain a finite area $\cal A_{\alpha}$ we include a time-independent component of the out-of-plane field, coming from a substrate or a back gate, such that $E^{z}=-\partial_{t}A^{z}+E^{z}_{0}$ (see Fig. 1c). In our simulations, we set $E_{1}=0.1$ V/nm and $E^{z}_{0}\approx 2/3E_{1}$, and tune $E_{2}=rE_{1}$. This electric field scale induces momentum displacements within the region of validity of the low-energy description, $eA_{j}\cos\theta_{j}\lesssim\gamma_{1}/2v_{F}$ [35].

### v2

@@ Abstract
Quantum systems coupled to light exhibit a range of unique emergent phenomena, providing a powerful platform to design and control phases of matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. It was recently appreciated that the paradigm of Floquet engineering can be generalized to multi-chromatic driving, leading to even richer physics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 31]. In this setting, the photon numbers of distinct driving modes act as synthetic lattice coordinates, opening access to higher-dimensional topological phases that support novel pumping mechanisms. A prime example is topological frequency conversion, whereby energy is transferred between two modes with frequencies $f_{1}$ and $f_{2}$ at a rate given by the power quantum $hf_{1}f_{2}$.
In the solid-state realm, Weyl semimetals were recently proposed as a platform to realize this effect, leveraging the sources and sinks of Berry curvature carried by their Weyl points [31]. In order to achieve optical amplification with topological frequency conversion in Weyl semimetals, electronic relaxation must be slow enough to ensure that the energy conversion rate exceeds that of dissipative loss. Ref. [31] estimates an amplification threshold for the relaxation time $\tau$ at $\sim 100~{}{\rm ps}$ for realistic material parameters—orders of magnitude larger than the sub-ps timescales typically found in experiments [33, 34]. These challenges motivate the search for alternative solid-state platforms for topological frequency conversion.
@@
Here we identify rhombohedral multilayer graphene (RMG) [35, 36] as a promising candidate towards this goal. Similarly to Weyl semimetals, RMG hosts Weyl nodes, but in a hybrid three-dimensional space formed by the two crystal momenta and the interlayer potential difference. This hybrid space can be dynamically addressed via the in-plane vector potential and out-of-plane electric field induced by the driving modes. Under suitable configurations of incident beams, a fraction of charge carriers in the system then converts energy between the modes at the quantized rate $hf_{1}f_{2}$, as illustrated in Fig. 1.
@@
Figure 1: (a) Rhombohedral multilayer graphene (RMG) can mediate topological frequency conversion between two incident driving modes with frequencies $f_{1}$, $f_{2}$. An interlayer potential $u$ gaps out the low-energy bands of RMG [shown for (b) bilayers and (c) trilayers], with a Berry curvature distribution [color scale] controlled by ${\rm sgn}(u)$. (d) The driving modes couple to electrons via their in-plane vector potential and perpendicular electric field, contained in the 3D vector $\bm{\alpha}$ [Eq. 3]. The Berry curvature ${\bm{\Omega}}$ in $\bm{\alpha}$ space (black arrows) hosts a monopole, or Weyl node, located at $(-k^{x},-k^{y},0)$ for an electron with wavevector $(k^{x},k^{y})$. Under quasi-periodic driving, the system explores a 2D surface ${\cal S}_{\bm{\alpha}}$, shown here for the drive configuration in a. Frequency conversion power is given by the quantized flux of Berry curvature $\mathbf{\Omega}$ through ${\cal S}_{\bm{\alpha}}$, see Eq. (5). (e) The intersection between ${\cal S}_{\bm{\alpha}}$ and the $E^{z}=0$ plane defines a momentum-space area $\cal A_{\bm{\alpha}}$ where electrons contribute to conversion power at a quantized rate ±$\pm hf_{1}f_{2}L$. Contours are labeled by the ratio of electric fields $E^{z}_{0}/E_{1}$ (see main text) and indicate the corresponding boundaries of $\cal A_{\bm{\alpha}}$.
The quasi two-dimensional nature of RMG furnishes key advantages for frequency conversion including reduced screening and lack of skin effects, along with gate tunability which provides an extra lever of control. Due to its pristine nature, relaxation rates in RMG are also expected to be much lower than in bulk materials, such as Weyl semimetals. Moreover, as a unique feature of our setup, the different couplings to in-plane and out-of-plane radiation fields lead to a strong sensitivity on the angle and polarization of the incident beams. Within the rhombohedral family, we find that Bernal bilayer graphene appears most promising for THz-scale implementations due to its lower propensity towards dissipation.
The two time-reversed valleys of RMG contribute to topological frequency conversion with an opposite sign; a nonzero net conversion rate hence requires an imbalance in the electronic distributions of the two valleys. We conclude our paper by discussing various strategies for inducing such valley-asymmetric electronic distributions.
@@
Setup. RMG consists of $L$ graphene sheets stacked such that each $A$ sublattice lies above the $B$ sublattice of the layer below—these superimposed sites are coupled by the leading interlayer tunneling term $\gamma_{1}\approx 360$ meV. Near charge neutrality, the wavefunctions of carriers primarily reside in the subspace ($A_{1},B_{L}$) comprising the leftover $A$ sites of the bottom layer and $B$ sites of the top layer. The system can thus be described by an effective two-level Hamiltonian in each of the two valleys [36], which in the presence of an out-of-plane electric field $E^{z}$ reads
$H_{\xi}(\mathbf{k})=\begin{pmatrix}u/2&-\gamma_{L}\left(\xi k^{x}-ik^{y}\right)^{L}\\ -\gamma_{L}\left(\xi k^{x}+ik^{y}\right)^{L}&-u/2\end{pmatrix},$ (1)
@@
with the wavevector $\mathbf{k}=(k^{x},k^{y})$ measured from the hexagonal Brillouin zone corner corresponding to valley $\xi=\pm 1$. Here $\gamma_{L}=(\hbar v_{F})^{L}/\gamma_{1}^{L-1}$, with $v_{\rm F}=8.4\times 10^{5}$ m/s the Fermi velocity of graphene, while $u=d(L-1)eE^{z}$ denotes the potential difference induced by $E^{z}$, with $d=0.33$ nm the interlayer distance and $e$ the electron charge. For $u=0$ the low-energy excitations of RMG are chiral fermions with an $L$-th power dispersion [37]. A nonzero interlayer potential $u$ opens a gap in the dispersion and controls the Berry curvature of the low-energy bands (Fig. 1b,c)—a feature at the heart of our proposal.
@@
Bichromatic driving induces an electric field ${\bf{E}}(t)={\bf{E}}_{1}(t)+{\bf{E}}_{2}(t+\Delta t)$ in the system. Here ${\bf{E}}_{j}(t)$ is the monochromatic field from mode $j$ with frequency $f_{j}$, and $\Delta t$ controls the relative phase shift between the modes—relevant when $f_{1}$ and $f_{2}$ are commensurate. Due to the quasi-2D nature of RMG we disregard skin effects and plasmon excitations that can become important in 3D implementations such as Weyl semimetals [31]. We further consider THz-scale frequencies, corresponding to wavelengths $\lambda\sim 100\,\rm\mu m$, such that the electric field can be considered uniform in the sample. The in-plane electric field component ${\bf{E}}^{\parallel}$ couples to electronic crystal momenta via the Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}+e\mathbf{A}^{\parallel}(t)/\hbar$, with ${\bf{A}}^{\parallel}(t)$ the net vector potential induced by ${\bf{E}}^{\parallel}(t)$. In contrast, the out-of-plane component $E^{z}$ couples through the effective dipole moment in the low-energy subspace, Eq. (1). The coupling of radiation thus strongly depends on the incidence angles $\theta_{j}$ and polarization of the drives (see Fig. 1a), a unique feature of this setup.
@@
${\bf{d}}_{\xi}(\mathbf{k},t)=\gamma_{L}\left(-\text{Re}[\Pi_{\xi}^{L}(\mathbf{k},t)],-\text{Im}[\Pi_{\xi}^{L}(\mathbf{k},t)],\alpha^{z}(t)\right).$ (2)
@@
Below, we find it convenient to recast the time dependence of $\bm{\alpha}_{j}$ (and therefore $\mathbf{d}_{\xi}$) in terms of the phases of the modes, $\phi_{j}=\omega_{j}t$ with $\omega_{j}=2\pi f_{j}$.
@@
When the time dependence of $H_{\xi}(\mathbf{k},t)$ is quasi-adiabatic, $P_{\rm FC}^{\xi}(\mathbf{k})$ is proportional to the number of times the surface spanned by the unit vector $\hat{\mathbf{d}}_{\xi}\equiv{\bf{\mathbf{d}_{\xi}}}/|{\bf{\mathbf{d}_{\xi}}}|$ encloses the origin [17, 31, 38],
$P_{\rm FC}^{\xi}(\mathbf{k})=-hf_{1}f_{2}\int_{0}^{2\pi}\frac{d\phi_{1}d\phi_{2}}{4\pi}\hat{{\bf{d}}}_{\xi}\cdot(\partial_{\phi_{1}}\hat{{\bf{d}}}_{\xi}\times\partial_{\phi_{2}}\hat{{\bf{d}}}_{\xi}).$ (4)
This result is derived in the Appendix and emerges from the leading-order diabatic correction to the dynamics—along with the fact that, for incommensurate frequencies, the system uniformly explores the two-dimensional parameter space $0\leq\phi_{j}\leq 2\pi$. For commensurate frequencies, this formula is recovered by also averaging $P_{\rm FC}^{\xi}(\mathbf{k})$ over $\Delta t$. Changing variables from $\phi_{j}$ to $\bm{\alpha}$ yields
@@
In analogy to Weyl semimetals, sources and sinks of Berry curvature in the hybrid three-dimensional space $\bm{\alpha}$ are quantized and confined to topologically-protected gap-closing points, analogous to Weyl nodes  [^1] The term Weyl node normally refers to nodal points in the 3D parameter space comprising the crystal momenta in bulk materials. Here we use the term “Weyl node” in a generalized sense to describe gap closings as a function of any three parameters.. Indeed, the RMG Hamiltonian in Eq. (2) hosts a charge-$L$ Weyl node at $\bm{\alpha}_{0}(\mathbf{k})=(-k^{x},-k^{y},0)$,
$\nabla_{\bm{\alpha}}\cdot{\mathbf{\Omega}_{\xi}}(\mathbf{k},\bm{\alpha})=2\pi\xi L\delta(\bm{\alpha}-\bm{\alpha}_{0}(\mathbf{k})).$ (6)
A valence band electron with wavevector $\mathbf{k}=(k^{x},k^{y})$ therefore acts as a topological frequency converter if $\bm{\alpha}_{0}(\mathbf{k})$ is located within the surface $\cal S_{\bm{\alpha}}$. This condition defines an area $\cal A_{\bm{\alpha}}$ in the Brillouin zone (see Fig. 1e) where each valence-band electron with momentum $\mathbf{k}\in\cal A_{\bm{\alpha}}$ mediates energy transfer at the quantized rate $P_{\rm FC}^{\xi}(\mathbf{k})=\pm hf_{1}f_{2}\xi L$. Here, the $+$ and -- signs occur when the area element $d{\bf{S}}_{\bm{\alpha}}$ defined from the driving protocol [below Eq. (5)] points inwards and outwards, respectively. For electrons outside of $\cal A_{\bm{\alpha}}$, $P_{\rm FC}^{\xi}(\mathbf{k})=0$.
We now consider the dynamics of the many-body state. In the regime of adiabatic driving and slow relaxation relative to the driving frequencies (that is, $\omega_{j}\tau\gg 1$), the steady-state population of band $n=0,1$ near valley $\xi$ takes an approximately stationary value, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$  [^2] In a simple relaxation-time approximation, $\bar{\rho}_{n}^{\xi}(\mathbf{k})$ is given by the time-averaged value of $f_{\beta}([\varepsilon_{n}(\mathbf{k},t)-\mu])$ with $f_{\beta}(E)$ the Fermi-Dirac distribution [31]; we expect more general dissipation mechanisms will lead to qualitatively similar distributions.. The frequency conversion intensity (i.e. the conversion power per unit area) from valley $\xi$ is obtained by weighing the individual contribution $P_{\rm FC}^{\xi}(\mathbf{k})$ of each electron with the corresponding steady-state occupation,
${\cal I}_{\rm FC}^{\xi}=\pm hf_{1}f_{2}L\xi\int_{\cal A_{\bm{\alpha}}}\frac{{\rm d}^{2}\mathbf{k}}{(2\pi)^{2}}\left(\bar{\rho}_{0}^{\xi}(\mathbf{k})-\bar{\rho}_{1}^{\xi}(\mathbf{k})\right).$ (7)
The boundary of $\cal A_{\bm{\alpha}}$ consists of the wavevectors $\mathbf{k}$ for which the instantaneous gap of $H_{\xi}(\mathbf{k},t)$ closes at some time $t$. States near this boundary thus undergo inter-band Landau-Zener tunneling, which causes significant dissipation through Landau-Zener absorption—whereby electrons pumped to the valence band relax dissipatively, leading to an irreversible loss of the invested energy.
@@
where we used $A_{j}\sim E_{j}/\omega_{j}$. For electric fields $E_{j}\sim 0.1$ V/nm and $\theta_{j}\sim 89^{\circ}$, ${\cal I}_{\rm FC}^{\xi}\sim 10L$ W/mm2 (see also numerics in Fig. 3). Interestingly, the conversion intensity is proportional to the geometric mean of radiation intensities, $I=c\epsilon_{0}E_{1}E_{2}/2$, through ${\cal I}_{\rm FC}^{\xi}/I\approx 4L\alpha_{\rm EM}\cos^{2}(\theta)/\pi^{2}$ with $\alpha_{\rm EM}$ the fine-structure constant.
When the chemical potential $\mu_{\xi}$ in a given valley moves away from charge neutrality, we expect the conversion intensity $I_{\rm FC}^{\xi}$ to decrease, due to cancellations between conduction and valence-band carriers (for $\mu_{\xi}>0$) or de-populating the valence band (for $\mu_{\xi}<0$). Because of the opposite contributions from carriers in valleys $\xi=\pm 1$ (a consequence of time-reversal symmetry), a net effect requires a valley imbalance in the number of active frequency converters, e.g. through unequal chemical potentials, $\mu_{+}\neq\mu_{-}$, or a valley-selective population inversion. We discuss strategies to induce such imbalances below.
Numerical simulations. We support our conclusions with simulations of Bernal bilayer graphene subjected to bichromatic radiation, including trigonal warping corrections to its low-energy band structure [41]. Following Ref. [31], we consider the master equation
$\partial_{t}\hat{\rho}_{\xi}(\mathbf{k},t)=-\frac{i}{\hbar}[\hat{H}_{\xi}(\mathbf{k},t),\hat{\rho}_{\xi}(\mathbf{k},t)]+{\cal D}_{\xi}(\mathbf{k},t)\left\{\hat{\rho}_{\xi}(\mathbf{k},t)\right\},$ (9)
where $\hat{H}_{\xi}(\mathbf{k},t)$ and $\hat{\rho}_{\xi}(\mathbf{k},t)$ denote the second-quantized Hamiltonian and density matrix of the system. We model dissipation via ${\cal D}_{\xi}(\mathbf{k},t)\left\{\hat{\rho}_{\xi}(\mathbf{k},t)\right\}=-\frac{1}{\tau}(\hat{\rho}_{\xi}(\mathbf{k},t)-\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t))$, characterizing uniform relaxation on a timescale $\tau$ towards an instantaneous equilibrium state with a bath at inverse temperature $\beta$, $\hat{\rho}_{\xi}^{\rm eq}(\mathbf{k},t)=e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}/{\rm Tr}[e^{-\beta[\hat{H}_{\xi}(\mathbf{k},t)-\mu_{\xi}\hat{n}]}]$, with $\hat{n}$ the electron density operator. For simplicity we work at charge neutrality, $\mu_{\xi}=0$; we consider effects of driving in the presence of a Fermi surface (for $\mu_{\xi}\neq 0$) in the Appendix.
Figure 2: Momentum-resolved frequency conversion in Bernal bilayer graphene. We consider two linearly-polarized modes in the configuration of Fig. 1a with $\left(\pi/2-\theta_{j}\right)\approx 0.1^{\circ}$, electric field scale $E_{1}=0.1$ V/nm, relaxation time $\tau=500$ ps, and at charge neutrality, $\mu=0$. Contributions to frequency conversion from valleys $\xi=+1$ (a,d) and $-1$ (b,e) are shown alongside dissipated power (c,f). Frequencies $f_{1}\approx 0.128$ THz and $f_{2}=f_{1}/\sqrt{2}$ [(a)–(c)] or $f_{2}=2f_{1}/3$ [(d)–(f)] illustrate quasi-periodic and periodic driving. The color scales represent time-averaged powers in units of $hf_{1}f_{2}$. The overlaid contour shows the predicted boundary of ${\cal A}_{\bm{\alpha}}$.
@@
This expression can be derived from first principles and, being agnostic to our theoretical analysis above, serves as an independent check of its conclusions.
@@
We consider two linearly-polarized modes propagating with incidence angles $\theta_{j}$ (relative to $\hat{\mathbf{z}}$) in the $xz$ and $yz$ planes, as depicted in Fig. 1a. The induced vector potentials are given by $\mathbf{A}_{1}(t)=A_{1}(\cos\theta_{1}\cos[\omega_{1}t],0,\sin\theta_{1}\cos[\omega_{1}t])$ and $\mathbf{A}_{2}(t)=A_{2}(0,-\cos\theta_{2}\cos[\omega_{2}(t{+\Delta t})],\sin\theta_{2}\cos[\omega_{2}(t{+\Delta t})])$, with $\Delta t$ controlling the phase shift between the modes. We consider large angles of incidence $\theta\approx\pi/2$, to compensate for the weaker effective coupling between the radiation fields and $\sigma^{z}$ in Eq. 2 (as compared to the in-plane components of the radiation). In order to obtain a finite area $\cal A_{\bm{\alpha}}$ we include a time-independent component of the out-of-plane field (coming from, e.g., a substrate or a back gate), such that $E^{z}=-\partial_{t}A^{z}+E^{z}_{0}$ (see Fig. 1d,e). In our simulations, we consider $E_{1}=0.1$ V/nm and $E^{z}_{0}\approx 2/3E_{1}$, and set $E_{2}=(\omega_{2}/\omega_{1})E_{1}$ so that the conversion region $\cal A_{\bm{\alpha}}$ remains unchanged for different frequency ratios. This electric field scale induces momentum displacements within the region of validity of the low-energy description, $eA_{j}\cos\theta_{j}\lesssim\gamma_{1}/2v_{F}$ [37].